1. Please solve the following by factoring:
a) x^2- 6x - 16 = 0
b) 6x^2 + 42x = 0
2. If f(t) = 2t^2 - 4t - 1, find
a) f(2)
b) f(-1)
Please help, I need step by step instructions.

Answer by stanbon(75887)

1. Please solve the following by factoring:
a) x^2- 6x - 16 = 0
(x-8)(x+2)=0
x=8 or x=-2
--------------------
b) 6x^2 + 42x = 0
6x(x+7)=0
x = 0 or x=-7
--------------------
2. If f(t) = 2t^2 - 4t - 1, find
a) f(2) = 2(2)^2-4*2-1 = -1
b) f(-1)= 2(-1)^2 -4*-1 -1 = 2+4-1=5
=================
Cheers,
Stan H.